OpenStudy (anonymous):

I need help on this question.-32/45 *48/35

OpenStudy (anonymous):

can some one help me

OpenStudy (anonymous):

since its multiplication you can just multiply the numerators by each other and then the denominators by each other. next step is simplify.\[-32* 48= -1536\] (numerator) the denominator is found by \[45*35=1575\] so the answer is \[-1536/1575\] however it can simplify to -512/525 by dividing the numerator and denominator by 3! hope this helpss :)

OpenStudy (anonymous):

Wow you are really smart thank you

OpenStudy (anonymous):

what about this....7 7/9 *1 11/45 =???

OpenStudy (anonymous):

can you help me please?

OpenStudy (anonymous):

hello............:(